what is the volume of 7.53g of oxygen?

1
Expert's answer
2022-02-18T14:09:03-0500

Oxygen is a molecule thud RFM IS 16×2=32

MOLES =7.53/32=0.2353125

AT RTP VOLUME OF GAS IS 24 SO 0.2353125×24=5.6475

=5.6475dm^3